What's the impulse of a *thing* if it's mass is 1kg and it's kinetic energy is 8J?



1
Expert's answer
2021-02-17T10:56:09-0500

Let's first find the velocity of the *thing* from the definition of the kinetic energy:


"KE=\dfrac{1}{2}mv^2,""v=\sqrt{\dfrac{2KE}{m}}."

Finally, we can find the impulse of the *thing*:


"J=\Delta p=mv=m\sqrt{\dfrac{2KE}{m}}=\sqrt{2mKE},""J=\sqrt{2mKE}=\sqrt{2\cdot1\ kg\cdot8\ J}=4\ N\cdot s."